The probability that Janice smokes is 3/10. The probability that she Smokes and develops lung cancer is 4/15. Find the probability that Janice develops lung cancer, given that she smokes.
1
Expert's answer
2019-07-22T03:51:57-0400

A ='Janice smokes',P(A) = 3/10

B=' develops lung cancer',P(B and A) = 4/15,

P(B|A) = P(B and A) / P(A) = 4/15 * 10/3 = 8/9.